# Ionic association and Wien effect in 2D confined electrolytes

### v1

@@ Introduction
Recent years have witnessed critical advances in the fabrication of atomic-scale fluidic channelsBocquet and Charlaix (2010); Garaj et al. (2010); Lee et al. (2010); Feng et al. (2016); Secchi et al. (2016); Radha et al. (2016); Esfandiar et al. (2017), in an effort to achieve molecular control over water and ion transport – similar to the transport machinery of cells. Considering the intense complexity of such nano-fabrication techniques, theoretical guidance has become necessary; however, traditional continuous models of fluid and charge transport breakdown under such extreme confinementBocquet and Charlaix (2010); Kavokine, Netz, and Bocquet (2021); Kavokine et al. (2019); Kavokine, Robin, and Bocquet (2022); Robin and Bocquet (2023); Richards et al. (2012); Robin et al. (2022); Robin, Kavokine, and Bocquet (2021).
@@
Particular attention has been drawn to 2D nanochannels recentlyRadha et al. (2016); Emmerich et al. (2022) (Figure 1(a)). These systems, which are fabricated by van der Waals assembly, consist in flakes of a 2D material (graphene, MoS2, hBN…) separated by graphene ribbons, creating an atomically-smooth and -thin channel permeable to water. Electrolytes confined inside such structures are expected to form a single ionic layer, which has been shown to enhance electrostatic interactions between ionsKavokine, Robin, and Bocquet (2022); Robin, Kavokine, and Bocquet (2021) and promote memory effects in conduction, such as the recently-demonstrated memristor effectRobin et al. (2022); Robin, Kavokine, and Bocquet (2021).
@@
The properties of ions in such systems have been studied using various techniques, both in theory – mean field approachesRobin, Kavokine, and Bocquet (2021); Levin (2002), exact field theoriesMinnhagen (1987); Robin et al. (2023)) –, and in simulations – BrownianRobin, Kavokine, and Bocquet (2021); Kavokine, Robin, and Bocquet (2022), molecularRobin, Kavokine, and Bocquet (2021); Zhao et al. (2021) or ab initio dynamicsZhao et al. (2021), and more recently machine learning forcefields trained with DFT-generated datasetsFong et al. (2024). Overall, a growing body of evidence is shedding light on links between electrostatic correlations and non-linear ion transport Avni et al. (2022). A common blind spot of the aforementioned numerical approaches is the description of out of equilibrium ion transport. While theoretical predictions recently showed that 2D nanochannels should display non linear conduction due to electrostatic interactions between ions, numerical evidence for this has remained scarce, in particular due to the high computational cost of simulating large systems over long timescales.
@@
In this work, we implement molecular dynamics simulations to carry out a full description of ion association under nanometric confinement, and show how electrostatic interactions impact charge transport at the molecular scale. We use a combination of Brownian dynamics (BD), where the solvent and channel walls are treated implcitly, and all-atom molecular dynamics (MD). The latter is particularly suited for identifying the impact of the discreteness of water molecules, while the former gives access to long-time dynamics by saving computational cost. We show that these two techniques offer complimentary tools to characterize ion association and ion-ion correlations in general. In both cases, we fully describe the formation of ionic clusters as function of the relative strength of electrostatic interactions compared to thermal noise.
@@
This paper is organized as follows. In Section I, we describe our numerical implementation. Section II describes ionic association. We show how correlation functions can be used to determine the fraction of paired ions at thermal equilibrium, unveiling a transition between a fully paired state and a partially dissociated state as function of temperature, akin a modified Kosterlitz-Thouless transition. In section III, we explore out-of-equilibrium properties and demonstrate non-linear transport of ions in the monolayer, governed by pair breaking under an electric field (a phenomenon known as the second Wien effect).
@@ I Methods
Two different simulation methods are used throughout this paper. In all-atom molecular dynamics (MD), the entire system (water molecules, ions and atoms from the channel walls, see Figure 1(a)) is simulated using classical Lennard-Jones and Coulombic forcefields and solving Newton’s equations of motion. In Brownian dynamics (BD), on the other hand, water and channel walls are treated implicitly, and only the positions of ions are tracked using overdamped Langevin equations. They are carried out using GROMACSAbraham et al. (2024) and LAMMPSThompson et al. (2022) softwares, respectively. Both equilibrium (EMD and EBD, without any external electric field) and non equilibrium simulations (NEMD and NEBD, with an external electric field) are used. The trajectory are obtained from simulation results using the MDAnalysisGowers et al. (2016) Python library, and visualized with OvitoStukowski (2010). Structure analysis of ion clusters is done using the network analysis Python library NetworkX Hagberg, Swart, and Schult (2008).
I.1 All-atom molecular dynamics
Molecular dynamics simulations consist of two sheets of fixed carbon atoms, with periodic boundary condition. In order to reproduce the crystalline structure of graphene, a rectangular unit cell of size $\sqrt{3}d_{c}\times 3d_{c}$ is created, corresponding to 4 carbon atoms, where $d_{c}=1.42\ \text{\AA}$ is the carbon-carbon distance (see Figure 1 for the definition of the unit cell). This cell is replicated 81 times along the $x$ direction and 46 times along the $y$ directions for each sheet. Both sheets are separated by a distance $h$, which creates a slit of height $h=7\ \text{\AA}$. Because of periodicity, we extend the simulation box to 20nm in the z direction to avoid interaction between periodically replicated images. This gives a total simulation box of size $19.9\ \text{nm}\times 19.6\ \text{nm}\times 20\ \text{nm}$. The inside of the slit is filled with water ad ions consisting in 2 identical mono-atomic species with a varying charge ±$\pm Ze$. The number of water molecules and ions are respectively $N_{w}=4500$ and $N_{+}=N_{-}=N=100$. This corresponds to a ionic concentration inside the slit of 0.6 M. The procedure used to estimate the number of water molecules in this geometry is presented in supplementary material (Section S2).
Water molecules are modeled using the SPC/E modelBerendsen, Grigera, and Straatsma (1987) (3 point charges model) and maintained rigid with the SHAKE algorithmRyckaert, Ciccotti, and Berendsen (1977). Short-range interactions are modeled using Lennard-Jones (LJ) potentials. The cut-off distance for LJ and Coulombic interactions is 1.2 nm. The atom mass, charges and Lennard Jones parameters are summarized in the supplementary material (Table S1). For both ions we use the same LJ parameters and masses (which are taken from the sodium force-field) with opposite charges. Long-range Coulombic interactions are treated with the smooth particle-mesh Ewald methodDarden, York, and Pedersen (1993); Essmann et al. (1995). When initializing the simulation, the total energy is minimized using the steepest descent minimization algorithm.
Simulations are performed in the NVT ensemble using the velocity VerletSwope et al. (1982) algorithm with a timestep $\text{d}t=1$ fs. The system is coupled to a thermostat at 300K using velocity rescaling with a stochastic termBussi, Donadio, and Parrinello (2007). The simulation runs for $1.5\times 10^{7}\ \text{d}t=$15 ns, and the first 5 ns are discarded for thermalization during analysis.
I.2 Brownian dynamics
I.2.1 Equation of motion
In BD, ions evolve in a 2D continuous space representing the nanochannel. The following equation of motion is solved for each ion $i$:
$\frac{\text{d}\vec{r}_{i}}{\text{d}t}=\frac{D}{k_{B}T}\left(z_{i}eE\vec{u}_{x}-\vec{\nabla}U_{i}\right)+\sqrt{2D}\vec{\eta}_{i}(t),$ (1)
where $z_{i}=Z$ for cation and $z_{i}=-Z$ for anions, $E$ is the external field, $D$ is the diffusion coefficient (identical values for both cations and anions), $U_{i}$ is the total potential felt by the ion i, and the $\vec{\eta}_{i}$ are independent 2D white noise forces:
$\langle\vec{\eta}_{i}(t)\rangle=\vec{0},\quad\langle\vec{\eta}_{i}(t)\cdot\vec{\eta}_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$ (2)
The cut-off distance for the interaction between the ions is $r_{c}=20$ nm. The shape of the inter-ions potential is the object of the next subsection.
The strength of the interaction can be tuned either by changing the charge of the ions, or the temperature. Without explicit solvent, both techniques should yield identical results. In the following, we choose to vary the temperature. The diffusion coefficient is adjusted in order to keep the relaxation time constant, with value:
$\tau=\frac{Dm}{k_{B}T}=9.6\ \text{fs}$ (3)
We choose $m=23\ \text{uma}$ for both ions. This assures that the ionic mobility remains the same across simulations at different temperatures.
The system consists of $N_{+}=N_{-}=N$ ions in a 2D slit of dimension $L\times w$. In section II, we use $N=100$ ions, in a box of lendth and width $L=w=100$ nm, for a total simulations times of $10^{6}$ timesteps. In section III, we use $N=4000$ ions, in a box of length $L=2000$  nm and width $w=200$  nm, for a total simulations times of $10^{5}$ timesteps. For all BD simulations, the width $w$ of the slit is chosen to have a constant ionic density of $10^{16}\ \text{atom}.\text{m}^{-2}$ for each ions, corresponding to a concentration of 0.025 M. The timestep is set to 5 ps.
I.2.2 Interaction potential
In BD, water and channel walls are treated implicitly. In bulk simulations with no walls, this amounts to introducing the relative dielectric constant $\varepsilon_{w}$ of water; here, we also must take into account the dielectric properties of the walls, which in general strongly contrast with waterSchlaich, Knapp, and Netz (2016); Fumagalli et al. (2018). Dielectric mismatch between water and walls results can be accounted for through image charges within the wall in the presence of ions. Reference Kavokine, Robin, and Bocquet (2022) shows that in a slit of height $h$, with wall of permittivity $\varepsilon_{s}$ and water of permittivity $\varepsilon_{w}$, this effect renormalises ion-ion interaction, so that the pairwise potential becomes:
$\beta V_{ij}(r_{ij})=-\frac{\text{sign}(z_{i}z_{j})}{T^{*}}\log\left(\frac{r_{ij}+C}{r_{ij}+\xi}\right),$ (4)
where $z_{i}=\pm 1$ is the charge of the ion $i$, $\xi$ is the dielectric length, $T^{*}$ is the reduced Coulomb temperature, a dimensionless measure of temperature, $C\,(\sim\text{\AA})$ is a short distance cut-off, $r_{ij}$ is the distance between the ions and $\beta=1/k_{B}T$.
The dielectric length $\xi$ depends only on the geometry of the slit, and not on the properties of the ions. We have:
$\xi=\frac{\varepsilon_{w}}{2\varepsilon_{s}}h=13.8\text{nm},$ (5)
for ions in water ($\varepsilon_{w}\sim 80$) in a slit of height $h=7\text{\AA}$, confined by walls with a very low dielectric constant ($\varepsilon_{s}=2$). This length corresponds to the typical length below which the electric field created by the ions is confined because of dielectric contrast. While in MD simulations electrostatic interactions follow the Coulomb’s law, it is useful to notice that they correspond to $\varepsilon=1$ (or $\xi=27.6$ nm), since we model channel walls by a single carbon atomic layer followed by a slab of empty space. An other interesting limit is the case of a total confinement of the electric field $\xi\to\infty$, for which we recover the logarithmic potential of a 2D Coulomb gas.
The short range cutoff $C=3\text{\AA}$ is an additional parameter that is chosen to avoid divergences of the potential at short ionic distances, without requiring an additional repulsive potential.
The reduced Coulomb temperature $T^{*}$ is an non dimensional parameter which measures the inverse of the interaction strength:
$T^{*}=\frac{2\pi\varepsilon_{0}\varepsilon_{w}h}{Z^{2}e^{2}\mathcal{K}}k_{B}T,$ (6)
where $\mathcal{K}=1.11$ is a geometric factor Robin, Kavokine, and Bocquet (2021). The interaction can be controlled by changing either the charges of the ions, or the temperature. In relevant experimental conditions, $T$ is usually fixed to 300K, and $Z$ is varied by using salts of various valence, $T^{*}=0.1$ corresponds to divalent ions and $T^{*}=0.4$ to monovalent ions. For practical reason, we rather fix in this work $Z=2$ and vary the temperature $T$ as a way of tuning interactions continuously over a wide range of strengths.
Finally, the total effective interaction potential felt by an ion in Brownian dynamics simulation is given by:
$U_{i}=\sum_{j\neq i}V_{ij}^{c}(r_{ij}),$ (7)
where:
$V_{ij}^{c}(r_{ij})=\left\{\begin{array}[]{cc}V_{ij}(r_{ij})&\quad\text{if }r_{ij}<r_{c}\\ 0&\quad\text{otherwise}\end{array}\right.$ (8)
I.3 Convention for correlations functions
The total pair correlation function is defined as:
$g_{p}(\vec{r})=\frac{\mathcal{A}}{4N^{2}}\left\langle\sum_{i}^{2N}\sum_{j\neq i}^{2N}\delta(\vec{r}-(\vec{r}_{i}-\vec{r}_{j}))\right\rangle,$ (9)
[…]
$g_{p}(r)=\frac{1}{2\pi r}\frac{\mathcal{A}}{4N^{2}}\left\langle\sum_{i}^{2N}\sum_{j\neq i}^{2N}\delta(r-|\vec{r}_{i}-\vec{r}_{j}|)\right\rangle.$ (10)
We can alternatively define $g_{\text{op}}$ and $g_{\text{id}}$ the pair correlation functions for respectively oppositely charged and identically charged ions :
$\begin{array}[]{ccc}g_{\text{op}}(r)&=&\frac{1}{2\pi r}\frac{\mathcal{A}}{N^{2}}\left\langle\sum_{i,\text{cations}}^{N}\sum_{j,\text{anions}}^{N}\delta(r-|\vec{r}_{i}-\vec{r}_{j}|)\right\rangle\\ g_{\text{id}}(r)&=&\frac{1}{2\pi r}\frac{\mathcal{A}}{N^{2}}\left\langle\sum_{i,\text{cations}}^{N}\sum_{j\neq i,\text{cations}}^{N}\delta(r-|\vec{r}_{i}-\vec{r}_{j}|)\right\rangle\end{array}$ (11)
$g_{\text{op}}$ is invariant under inversions of anions and cations, however $g_{\text{id}}$ is slightly asymmetric - even for symmetric anion and cation - because of the structure of water molecules. In the following we take the average of both anion-anion and cation-cation pair correlation functions.
In practice, we compute all the $2N^{2}N_{t}$ unique ionic distances, where $N_{t}$ is the number of simulation steps used, and $N$ is the number of ions. The distance between ions are corrected to take into account the periodicity. Radial pair histograms are computed every 5 ns ($10^{3}$ timesteps), for a total simulations length of 5 $\mu$s ($10^{6}$ timesteps), and the first 500 ns are discarded. In order to estimate the uncertainty on those quantity, we split the time interval into 10 windows, and we compute the standard deviation of the window averages.
The running coordination number $N_{\text{op}}(r)$ and $N_{\text{id}}(r)$ can be obtained upon integration of the radial pair correlation function of respectively opposite and identical ions. It corresponds to the number of ions in a shell of radius $r$ around any ions:
$N_{\text{op}}(r)=2\pi\frac{N}{\mathcal{A}}\int_{0}^{r}g_{\text{op}}(u)u\text{d}u.$ (12)
$N_{\text{id}}(r)=2\pi\frac{N}{\mathcal{A}}\int_{0}^{r}g_{\text{id}}(u)u\text{d}u.$ (13)
The counter charge density around an ion can be defined as:
$q(r)=-\frac{1}{4\pi rN}\left\langle\sum_{i}^{2N}\sum_{j\neq i}^{2N}\text{sign}(z_{i}z_{j})\delta(r-|\vec{r}_{i}-\vec{r}_{j}|)\right\rangle,$ (14)
$q$ can be integrated to obtain:
$Q(r)=2\pi\int_{0}^{r}q(u)u\text{d}u=N_{\text{op}}(r)-N_{\text{id}}(r),$ (15)
which corresponds to the total counter charge in a disk of radius $r$ around an ion. Those functions are normalized by the central ion charge, such that $Q(0)=0$, and $Q(r\to\infty)=1$, which corresponds to global neutrality of the system.
Adding an external electric field breaks the isotropy of the system. In this case, the running coordination number $N(r,E)$ can still be defined, but it does not exactly have the same physical meaning, as the anisotropic correlation are integrated on a circle.
I.4 Ionic current
In simulations where an external field $E$ is applied, by using translation invariance along the direction of the field, we compute the ionic current using:

### v2

@@ Introduction
Recent years have witnessed critical advances in the fabrication of atomic-scale fluidic channelsBocquet and Charlaix (2010); Garaj et al. (2010); Lee et al. (2010); Feng et al. (2016); Secchi et al. (2016); Radha et al. (2016); Esfandiar et al. (2017), in an effort to achieve molecular control over water and ion transport – similar to the transport machinery of cells. Considering the intense complexity of such nano-fabrication techniques, theoretical guidance has become necessary; however, traditional continuous models of fluid and charge transport breakdown under such extreme confinementBocquet and Charlaix (2010); Kavokine, Netz, and Bocquet (2021); Kavokine et al. (2019); Kavokine, Robin, and Bocquet (2022); Robin and Bocquet (2023); Richards et al. (2012); Robin et al. (2023a); Robin, Kavokine, and Bocquet (2021).
@@
Particular attention has been drawn to 2D nanochannels recentlyRadha et al. (2016); Emmerich et al. (2022) (Figure 1(a)). These systems, which are fabricated by van der Waals assembly, consist in flakes of a 2D material (graphene, MoS2, hBN…) separated by graphene ribbons, creating an atomically-smooth and -thin channel permeable to water. Electrolytes confined inside such structures are expected to form a single ionic layer, which has been shown to enhance electrostatic interactions between ionsKavokine, Robin, and Bocquet (2022); Robin, Kavokine, and Bocquet (2021) and promote memory effects in conduction, such as the recently-demonstrated memristor effectRobin et al. (2023a); Robin, Kavokine, and Bocquet (2021).
@@
The properties of ions in such systems have been studied using various techniques, both in theory – mean field approachesRobin, Kavokine, and Bocquet (2021); Levin (2002), exact field theoriesMinnhagen (1987); Robin et al. (2023b)) –, and in simulations – BrownianRobin, Kavokine, and Bocquet (2021); Kavokine, Robin, and Bocquet (2022), molecularRobin, Kavokine, and Bocquet (2021); Zhao et al. (2021) or ab initio dynamicsZhao et al. (2021), and more recently machine learning forcefields trained with DFT-generated datasetsFong et al. (2024). Overall, a growing body of evidence is shedding light on links between electrostatic correlations and non-linear ion transport Avni et al. (2022). A common blind spot of the aforementioned numerical approaches is the description of out of equilibrium ion transport. While theoretical predictions recently showed that 2D nanochannels should display non linear conduction due to electrostatic interactions between ions, numerical evidence for this has remained scarce, in particular due to the high computational cost of simulating large systems over long timescales.
@@
In this work, we implement molecular dynamics simulations to carry out a full description of ion association in 2D nanochannels, and show how electrostatic interactions impact charge transport at the molecular scale. We use a combination of Brownian dynamics (BD), where the solvent and channel walls are treated implcitly, and all-atom molecular dynamics (MD). The latter is particularly suited for identifying the impact of the discreteness of water molecules, while the former gives access to long-time dynamics by saving computational cost. We show that these two techniques offer complimentary tools to characterize ion association and ion-ion correlations in general. In both cases, we fully describe the formation of ionic clusters as function of the relative strength of electrostatic interactions compared to thermal noise.
@@
This paper is organized as follows. In Section I, we describe our numerical implementation and how it can be used to quantify ionic pairing in confinement. Section II describes ionic association. We show how electrostatic correlations under confinement result in ionic pairing at thermal equilibrium, unveiling a transition between a fully paired state and a partially dissociated state as function of temperature, similar to a modified Kosterlitz-Thouless transition. In section III, we explore out-of-equilibrium properties and demonstrate non-linear transport of ions in the monolayer, governed by pair breaking under an electric field (a phenomenon known as the second Wien effect).
@@ I Methods
In this work, we model an electrolyte confined in a 2D nanochannel of height $h$, made of carbon walls (Figure 1a). We assume that the system is connected to macroscopic fluidic reservoirs where Ag/AgCl electrodes are located, imposing a voltage drop across the system and creating a static electric field $\vec{E}$.
I.1 MD and BD dynamics
@@
Two different simulation methods are used throughout this paper. In all-atom molecular dynamics (MD), the entire system (water molecules, ions and atoms from the channel walls, see Figure 1(a)) is simulated using classical Lennard-Jones and Coulombic forcefields and solving Newton’s equations of motion. In Brownian dynamics (BD), on the other hand, water and channel walls are treated implicitly, and only the positions of ions are tracked using overdamped Langevin equations (Figure 1b). This is motivated by the finding that electrostatic interactions on ions from the walls and the solvent can be accounted for in a coarse-grained way by renormalizing the ion-ion interaction potential Kavokine, Robin, and Bocquet (2022). Further details regarding the simulation methodology can be found in Supplementary Information (SI).
I.2 Interaction potentials
In MD dynamics, as stated above, all particles (ions, wall atoms and water molecules) interact both through electrostatic and Lennard-Jones interactions, the latter representing attractive and repulsive van der Waals forces. The interaction energy between particles $i$ and $j$ thus reads:
$V_{ij}(r)=4\epsilon\left[\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\right]+\frac{q_{i}q_{j}}{4\pi\epsilon_{0}r},$ (1)
where $q_{i}$ is the charge of particle $i$, $r$ the interparticle distance and $\epsilon$ and $\sigma$ some particle-specific constants detailed in SI.
One of the main shortcomings of MD dynamics is that only a fraction of the simulation time is spent on simulating ion dynamics – the focus of this work. Integrating out solvent degrees of freedom allows ton run much faster implicit-solvent BD simulations. This is usually done by modelling water as a dielectric medium of constant $\epsilon_{w}=80$; however, here one must accounts for atomic confinement as well. Walls not only restrict ion transport geometrically, but also modify the electric potentials generated by ions due to having a much lower permittivity compared to water (in what follows, we will assume that the walls have a dielectric constant $\epsilon_{s}=2$) – this process, called interaction confinement Kavokine, Robin, and Bocquet (2022), is caused by the fact electric field lines from ions cannot fully penetrate the walls and instead concentrate within the fluid. Overall, this results in a renormalized ion-ion potential given byKavokine, Robin, and Bocquet (2022):
$V_{ij}(r)=\frac{q_{i}q_{j}}{4\pi\epsilon_{0}\epsilon_{w}r}\int_{0}^{\infty}J_{0}(u)\frac{\tanh\left[u\frac{h}{2r}\right]+\frac{\epsilon_{w}}{\epsilon_{s}}}{\frac{\epsilon_{w}}{\epsilon_{s}}\tanh\left[u\frac{h}{2r}\right]+1},$ (2)
where $J_{0}$ is a Bessel function and we recall that $h$ is the height of the channel. For $\epsilon_{w}/\epsilon_{s}\gg 1$ and $r\gg h$, the above integrand can be cut into $0<u<A$ and $u>A$ (defined by $\int_{0}^{A}J_{0}(u)\,\mathrm{d}u=0$, $A\simeq 1.1$). Assuming that the latter part oscillates quickly and integrate to zero, and that $\int_{0}^{A}J_{0}(u)\times[...]\simeq\int_{0}^{A}\frac{1}{A}\times[...]$, we obtain:
$\beta V_{ij}(r)=-\frac{\text{sign}(q_{i}q_{j})}{T^{*}}\log\left(\frac{r}{r+\xi}\right),$ (3)
where we introduced the inverse temperature $\beta$ as well as two other parameters: the dielectric length:
$\xi=A\epsilon_{w}h/2\epsilon_{s},$ (4)
and the (dimensionless) reduced temperature $T^{*}$ defined by:
$T^{*}=\frac{2\pi\varepsilon_{0}\varepsilon_{w}h}{Z^{2}e^{2}A}k_{B}T,$ (5)
with $Z$ the valency of ions and $e$ the elementary charge. This parameter plays the role of a coupling constant; we will use it throughout the manuscript as a measure of the strength of electrostatic interactions. One finds that $T^{*}\simeq 0.4$ for monovalent ions at room temperature and $T^{*}\simeq 0.1$ for divalent ions. In practice, we use this potential in BD simulations, with $\xi\simeq 14\,$nm. It should be noted that, in the limit $r\ll\xi$, one recovers $V_{ij}(r)\simeq q_{i}q_{j}/4\pi\epsilon_{0}\epsilon_{s}r$ as field lines now fully permeate the solid.
When comparing with MD simulations, one should notice that the system is confined between two vacuum slabs (as walls are made of a single layer of carbon atoms), corresponding to $\epsilon_{s}=1$ and $\xi\simeq 28\,$nm. Throughout the paper, to simplify discussion, we always assume that $Z=2$ and we instead use temperature as a tuning parameter of the interactions.
Confinement has been shown to reduce the effective permittivity of water or make it anisotropicFumagalli et al. (2018); Schlaich, Knapp, and Netz (2016); Monet et al. (2021). The exact dielectric properties of confined water are still debated however, with some recent experiments showing an increase in in-plane permittivity insteadWang et al. (2024). It has been shown elsewhere Robin, Kavokine, and Bocquet (2021) that dielectric anisotropy only has a weak effect on ion-ion interactions. Besides, our results can be extended in a straightforward way to other values of $\epsilon_{w}$ in confinement by tuning $T^{*}$.
Lastly, while in this paper we only consider the case of non-conducting walls made of a dielectric material, the impact of wall conductivity has been discussed elsewhereKavokine, Robin, and Bocquet (2022); Rochester et al. (2013); Kondrat et al. (2011). Briefly, conduction electrons within the walls are able to screen off ion-ion electrostatic interactions, reducing their range.
I.3 Ionic current
For a given value of the imposed electric field $\vec{E}$, we extract the ionic current density from simulated trajectories through: